Find the vertex, the line of symmetry, the max or min value of the guadratic function, and the graph function
f(x)=4x^2-8x+6

Answer by TimothyLamb(4379)

f(x) = 4x^2 - 8x + 6
---
the above quadratic equation is in standard form, with a=4, b=-8, and c=6
---
to solve the quadratic equation, plug this:
4 -8 6
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
x-intercepts:
the roots of the quadratic are complex
1 + (0.707106781)i
1 - (0.707106781)i
therefore the equation does not have x-intercepts
---
y-intercept:
f(0) = 4(0)^2 - 8(0) + 6 = 6
( 0, 6 )
---
min or max vertex:
vertex is a minimum at ( 1, 2 )
---
line of symmetry:
x = 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php